# Authorship Drift: How Self-Efficacy and Trust Evolve During LLM-Assisted Writing

### v1

@@ 1. Introduction
Large language models (LLMs) have emerged as powerful general-purpose tools, leveraging their natural language capabilities and vast knowledge across diverse domains (Brown et al., 2020; Zhao et al., 2023). Hence, LLMs are now perceived as collaborative partners that co-produce content with users (Lee et al., 2022a), extending their individual capabilities.
@@ 2.2. User Interaction Patterns with AI Systems
However, as AI systems become increasingly capable at producing high-quality outputs (Zhao et al., 2023; Brown et al., 2020; Kasneci et al., 2023), outcome-oriented evaluation becomes less informative for understanding human-AI interaction (Lee et al., 2022b; Ibrahim et al., 2025). When AI can reliably generate good results, the critical question shifts from “what was produced” to “how did the collaboration unfold” and “what was the user’s experience” (Guo et al., 2025; Lee et al., 2022b). In this context, examining users’ interaction patterns becomes particularly important. Understanding these collaborative processes is essential because they reveal whether users maintain meaningful agency (Guo et al., 2025) and cognitive engagement in the interaction (Kosmyna et al., 2025), which directly impacts sustained learning (Zhai et al., 2024) and long-term growth (Kobiella et al., 2025).
To better understand these collaborative processes, researchers have begun examining user interaction patterns through analysis of prompting strategies and conversational exchanges. For instance, McNichols et al. (McNichols and Lan, 2025) analyzed student-AI interactions to understand how different prompting strategies shape learning outcomes and engagement patterns. Similarly, Mysore et al. (Mysore et al., 2025) examined in-the-wild interaction logs to identify prototypical human–AI collaboration behaviors, capturing recurring patterns in how users iteratively refine AI-generated text through follow-up prompts. Building on these approaches, our study investigates how users’ evolving internal states are reflected in their prompting strategies, in order to understand which prompting strategies and interaction patterns foster human agency in LLM-assisted writing.
@@ 2.3. Authorship in LLM-Assisted Writing
While previous research has shown that LLM-assisted writing improves authors’ productivity (Li et al., 2024b; Dhillon et al., 2024) and creative processes (Wan et al., 2024; Shaer et al., 2024; Reza et al., 2024), central to this collaboration is the problem of authorship (Huang et al., 2025; Hwang et al., 2025; He et al., 2025). As LLMs increasingly function as coauthors (Lee et al., 2022a), they blur the boundary between the contributions of humans and models, potentially diminishing human agency (Qin et al., 2025; Draxler et al., 2024) while complicating questions of ownership (Hwang et al., 2025). Hence, this raises fundamental questions about how to define, measure, and attribute authorship in the context of human-LLM collaborative writing.
@@ 3.1.1. Participants
We aimed to recruit a minimum of 300 participants after applying our exclusion criteria. This target was established following the criteria of VanVoorhis and Morgan (VanVoorhis et al., 2007), which recommend around 30 observations per cell to achieve 80% power in group comparisons. Our pilot study ($N=20$) indicated the least frequent trajectory pattern occurred in about 10% of the sample, thus a sample of 300 would provide approximately 30 participants in the smallest group, meeting power requirements.
@@ 5.1.2. Role of Prompting Strategies in Self-Efficacy Trajectories
Our findings show that users whose self-efficacy declined relied heavily on editing requests while those who recovered self-efficacy used more review requests (Section 4.3). This pattern can be understood through findings from the educational domain, showing that direct corrective feedback positions learners as passive recipients, while diagnostic feedback that identifies issues maintains learner agency and supports self-efficacy (Griffiths et al., 2023). These studies emphasize that effective feedback should provide users the ability to maintain control over revision decisions in determining when, where, and how to implement suggested changes.
@@ 5.2.2. Using Behavioral Signals to Examine Self-Efficacy Shifts
Moreover, how users integrate LLM outputs provides complementary insight into self-efficacy shifts. As self-efficacy declines, users tend to adopt a larger portion of the model’s suggestions (Section 4.4). Tracking how semantic similarity or lexical overlap between user-written text and model-generated text changes over time can reveal shifts from active authorship toward passive acceptance. Techniques such as keystroke logging (Leijten and Van Waes, 2013) enable capturing these incorporation dynamics, offering additional evidence of when users’ contributions begin to recede.
@@ 5.2.3. Opportunities for Supporting Authorship in LLM-Assisted Writing
One opportunity lies in helping users stay aware of their interaction patterns and shift toward more reflective engagements over time. Our results revealed that users with declining self-efficacy tended to fall into repetitive “editing loops” of requesting direct revisions (Section 4.3.2), while those who regained self-efficacy more often sought evaluation rather than relying on the model to fix their text (Section 4.3.1). Since prior work suggests that making people’s own thoughts visible can foster metacognitive awareness (Tankelevitch et al., 2024), systems could help users recognize their engagement patterns by making collaboration history more visible (Hoque et al., 2024) or steering responses toward review and reflection (Yuan and Hu, 2024). Such interventions may help users move from passive acceptance of suggestions toward more intentional collaboration, with greater agency over their writing.
@@ 6. Limitations and Future Work
Finally, this study did not account for individual or system-related factors that prior work has shown to influence self-efficacy and trust, such as task proficiency (Qian and Wexler, 2024) or variations in system tone and expression (Zhou et al., 2024; Kim et al., 2024). Our analysis also used a single LLM, so we could not examine how such stylistic and behavioral differences across models might alter user perceptions (Fanous et al., 2025). Future research could extend this work by studying more diverse user groups and varying LLM configurations to better understand how these contextual factors interact with the dynamics of self-efficacy and trust.

### v2

@@ 1. Introduction
Large language models (LLMs) have emerged as powerful general-purpose tools, leveraging their natural language capabilities and vast knowledge across diverse domains (Brown et al., 2020; Zhao et al., 2025). Hence, LLMs are now perceived as collaborative partners that co-produce content with users (Lee et al., 2022), extending their individual capabilities.
@@ 2.2. User Interaction Patterns with AI Systems
However, as AI systems become increasingly capable at producing high-quality outputs (Zhao et al., 2025; Brown et al., 2020; Kasneci et al., 2023), outcome-oriented evaluation becomes less informative for understanding human-AI interaction (Lee et al., 2024; Ibrahim et al., 2025). When AI can reliably generate good results, the critical question shifts from “what was produced” to “how did the collaboration unfold” and “what was the user’s experience” (Guo et al., 2025; Lee et al., 2024). In this context, examining users’ interaction patterns becomes particularly important. Understanding these collaborative processes is essential because they reveal whether users maintain meaningful agency (Guo et al., 2025) and cognitive engagement in the interaction (Kosmyna et al., 2025), which directly impacts sustained learning (Zhai et al., 2024) and long-term growth (Kobiella et al., 2025).
To better understand these collaborative processes, researchers have begun examining user interaction patterns through analysis of prompting strategies and conversational exchanges. For instance, McNichols et al. (McNichols et al., 2025) analyzed student-AI interactions to understand how different prompting strategies shape learning outcomes and engagement patterns. Similarly, Mysore et al. (Mysore et al., 2025) examined in-the-wild interaction logs to identify prototypical human–AI collaboration behaviors, capturing recurring patterns in how users iteratively refine AI-generated text through follow-up prompts. Building on these approaches, our study investigates how users’ evolving internal states are reflected in their prompting strategies, in order to understand which prompting strategies and interaction patterns foster human agency in LLM-assisted writing.
@@ 2.3. Authorship in LLM-Assisted Writing
While previous research has shown that LLM-assisted writing improves authors’ productivity (Li et al., 2024b; Dhillon et al., 2024) and creative processes (Wan et al., 2024; Shaer et al., 2024; Reza et al., 2024), central to this collaboration is the problem of authorship (Huang et al., 2025; Hwang et al., 2025; He et al., 2025). As LLMs increasingly function as coauthors (Lee et al., 2022), they blur the boundary between the contributions of humans and models, potentially diminishing human agency (Qin et al., 2025; Draxler et al., 2024) while complicating questions of ownership (Hwang et al., 2025). Hence, this raises fundamental questions about how to define, measure, and attribute authorship in the context of human-LLM collaborative writing.
@@ 3.1.1. Participants
We aimed to recruit a minimum of 300 participants after applying our exclusion criteria. This target was established following the criteria of VanVoorhis and Morgan (Wilson Van Voorhis and Morgan, 2007), which recommend around 30 observations per cell to achieve 80% power in group comparisons. Our pilot study ($N=20$) indicated the least frequent trajectory pattern occurred in about 10% of the sample, thus a sample of 300 would provide approximately 30 participants in the smallest group, meeting power requirements.
@@ 5.1.2. Role of Prompting Strategies in Self-Efficacy Trajectories
Our findings show that users whose self-efficacy declined relied heavily on editing requests while those who recovered self-efficacy used more review requests (Section 4.3). This pattern can be understood through findings from the educational domain, showing that direct corrective feedback positions learners as passive recipients, while diagnostic feedback that identifies issues maintains learner agency and supports self-efficacy (Mutoni Griffiths et al., 2023). These studies emphasize that effective feedback should provide users the ability to maintain control over revision decisions in determining when, where, and how to implement suggested changes.
@@ 5.2.2. Using Behavioral Signals to Examine Self-Efficacy Shifts
Moreover, how users integrate LLM outputs provides complementary insight into self-efficacy shifts. As self-efficacy declines, users tend to adopt a larger portion of the model’s suggestions (Section 4.4). Tracking how semantic similarity or lexical overlap between user-written text and model-generated text changes over time can reveal shifts from active authorship toward passive acceptance. Techniques such as keystroke logging (Leijten and Waes, 2013) enable capturing these incorporation dynamics, offering additional evidence of when users’ contributions begin to recede.
@@ 5.2.3. Opportunities for Supporting Authorship in LLM-Assisted Writing
One opportunity lies in helping users stay aware of their interaction patterns and shift toward more reflective engagements over time. Our results revealed that users with declining self-efficacy tended to fall into repetitive “editing loops” of requesting direct revisions (Section 4.3.2), while those who regained self-efficacy more often sought evaluation rather than relying on the model to fix their text (Section 4.3.1). Since prior work suggests that making people’s own thoughts visible can foster metacognitive awareness (Tankelevitch et al., 2024), systems could help users recognize their engagement patterns by making collaboration history more visible (Hoque et al., 2024) or steering responses toward review and reflection (Yuan and Hu, 2025). Such interventions may help users move from passive acceptance of suggestions toward more intentional collaboration, with greater agency over their writing.
@@ 6. Limitations and Future Work
Finally, this study did not account for individual or system-related factors that prior work has shown to influence self-efficacy and trust, such as task proficiency (Qian and Wexler, 2024) or variations in system tone and expression (Zhou et al., 2025; Kim et al., 2024). Our analysis also used a single LLM, so we could not examine how such stylistic and behavioral differences across models might alter user perceptions (Fanous et al., 2025). Future research could extend this work by studying more diverse user groups and varying LLM configurations to better understand how these contextual factors interact with the dynamics of self-efficacy and trust.